OpenStudy (anonymous):

Kat has 19 coins, all quarters and dimes, that are worth a total of $4. The system of equations that can be used to find the number of quarters, q, and the number of dimes, d, she has is shown. q + d=19 0.25q + 0.1d=4 How many quarters does she have? 4 5 14 15

pooja195 (pooja195):

since we are trying to figure out the number of quarters we need to cancel out the $ (d) -0.1(q + d=19) 0.25q + 0.1d=4 Opposites cancel -0.1q -0.1 d=1.9 0.25q + 0.1d=4

pooja195 (pooja195):

Can you finish from here?

OpenStudy (anonymous):

no i cant